# Discovering the underlying analytic structure within Standard Model constants using artificial intelligence

### v2

@@ 6 $\exp(x)$ and $\log(x)$ ($\log$ base of $e$ of $x$), $\cos(x)$, $\sin(x)$, $\tan(x)$, $\arctan(x)$
2.3 Validation
As a sanity check, we have validated the genetic programming approach using known relationships. Although these relationships are not discussed in the context of the SM or any theory, they were established in the past, and well suited for our purpose of testing the algorithm.
The first relationship is [10]:
@@ 2.3 Validation
The second relationship for our benchmarks is [11]:
@@
Our program detected Eqs. 2.1 and 2.2 expressions using special test runs. The obtained simplicity ranks were 30 and 20, respectively. The difference $m_{H}-\sqrt{m_{Z}\>m_{t}}$ = -0.310 is greater than the absolute uncertainty allowed of 0.149 on the Higgs mass shown in Table 1. When GP finds this equation, it is not considered a solution.
3 Results
The expressions with the lowest analytic ranks 6 and 7 are shown in Tables LABEL:tab:long6 and LABEL:tab:long7, respectively. The regularities appear rather trivial, which can be checked by hand. Table LABEL:tab:long10 of the Appendix shows the results with the rank 10. Note that the analytic expressions with ranks 8 and 9 do not exist because of the default construction of the ranking system. We intentionally retained repeated snippets involving the same variables, as they may help model builders identify interesting patterns. In total, about 530,000 analytic relations were obtained [12] up to the rank 80, with and without rescaling by the $\rho(770)$ meson mass.
@@ 3 Results
The tables also show the relative difference for each analytic snippet. It is defined as $\Delta D=100\%\times|y-y_{t}|/y_{t}$ (expressed as a percentage), where $y_{t}$ is the target value and $y$ is the predicted value. All relations have $\Delta D<\varepsilon^{rel}$ on the target constants according to the definition of GP. The formulas with $\Delta D$ above 1% are not shown, to reduce the sizes of the tables. This requirement truncates the presentation for some targets, which usually correspond to $u,d$-masses, the CKM mixing angles and CP-violating phase, since they have relative uncertainties $\varepsilon^{rel}$ above 1%, see Table 1. Other target constants should not be affected by the 1% requirement, since their $\Delta D$ values are smaller than $\varepsilon^{rel}$ listed in Table 1.
The results shown in Tables LABEL:tab:long6 - LABEL:tab:long10 are dominated by the expressions for $\delta$, which has the largest absolute and relative uncertainties ($\varepsilon$ and $\varepsilon^{rel}$) among all the input variables. Among all the ranks obtained, leptons had the fewest solutions due to their relatively small mass uncertainty. All relationships from the symbolic regression have been preserved, even if they may appear highly unlikely. We remind that this choice reflects our intent to remain as theory-agnostic as possible, in case these analytic snippets are combined and simplified, and later used to inform or construct a model that connects all SM constants.
@@
Approximately 10% of the obtained relations satisfy the dimensional analysis if the rescaled masses listed in Table 1 are replaced with their original values expressed in physical units (MeV or GeV). Note that this replacement presumes the same numerical precision as that used for the rescaled masses in the GP inputs; otherwise, the expressions may not satisfy the GP output precision criteria.
@@
A preliminary visual inspection of the lowest-rank expressions reveals relationships that appear to be both remarkably simple and precise, warranting further investigation. For example, Table LABEL:tab:long10 has an expression $m_{t}=m_{W}+m_{W}\>\delta$ connecting the top-quark mass $m_{t}$ with the $W$ mass and $\delta$. The rank 15 snippets [^3] Tables with analytic ranks above 10 are not shown to reduce the size of the paper. contain another relation $m_{H}=m_{t}-(m_{b}+m_{t}/4)$ connecting the Higgs mass $m_{H}$ with $m_{t}$ and the $b-$quark mass. Another relation of the same rank is $m_{u}=2m_{e}/\sqrt{\theta_{12}}$, where $m_{u}$ is the mass of the $u$-quark, $m_{e}$ is the mass of the electron and $\theta_{12}$ is the CKM 12 mixing angle. The found expression $m_{H}=5\>m_{W}-3\>(m_{t}-m_{W})$ connects the Higgs mass with the top-quark mass and the $W$ mass. All such expressions are valid for the original mass units, without scaling by the $\rho(770)$ mass. The precision of these analytic relations significantly exceeds that of Eq. 2.2.
@@
Many relations involving the Higg mass are valid only for the dimensionless constants listed in Table 1. For example, $m_{H}=m_{t}-6\>\sqrt{m_{W}}$ connects $m_{H}$, $m_{W}$ and $m_{t}$. It does not pass the dimensional analysis when using the original masses, i.e. without rescaling the masses by the $\rho(770)$ meson mass. However, the mass of this meson can easily be substituted using Table LABEL:tab:long6 or LABEL:tab:long7, as we will discuss this later.
As an additional test, the GP studies were also conducted on the SM inputs without rescaling the lepton, quark, and heavy boson masses [1] by the $\rho(770)$ mass. In this case, the majority of snippets do not pass dimensional analysis, and the expressions for the electron and muon masses cannot be recovered due to the large precision on such masses. It was determined that the most interesting relations that pass dimensional analysis can also be recovered in the original analysis with the rescaled masses.
@@
One of the known property of the SM is that the masses of fermions, as well as the $m_{Z}$ and $m_{W}$ boson masses, tend toward zero when the Higgs boson mass is set to zero. This behavior arises from their dependence on the Higgs vacuum expectation value through the Yukawa couplings, which are free parameters in the theory [13]. We identified several simplest relations connecting the quark and $Z/W$ masses with the Higgs mass, which have the above mentioning property, and also pass the dimensional analysis. Here are some examples with small analytic ranks:
@@
These relations hold for the masses presented in Table 1, as well as for the original masses of quarks and bosons [1] with the physical units. Analytical expressions for lepton masses as linear functions of the Higgs mass were not found. Nevertheless, these examples illustrate that the reduction of the free parameters of the SM is possible, since 8 masses are expressed via the Higgs mass and 4 measured constants ($\alpha^{-1}$, $\delta$, $\theta_{12}$ and $\theta_{13}$). The origin of these expressions may be purely coincidental, and can only be understood within the framework of a specific model, which is beyond the scope of this paper.
@@
It will be interesting to assess the accuracy of the derived expressions as increasingly precise measurements of the SM parameters become available over time. The published relationships can be used for developing theories or as inputs for various artificial intelligence techniques to discover hidden analytic structures. For example, the analytic snippets can be simplified, combined, variables can be substituted, and a dimensional analysis can be applied after converting the masses back to their original physics units.
We should recall that the introduction of the non-fundamental particle in our study, the $\rho(770)$ meson, was made for convenience, to create the dimensionless inputs for GP. It is a parameter that defines the limits of numerical precision for the GP and, consequently, the applicability of any potential theory connecting the SM constants, should such a theory exist. In principle, any particle can be used for this purpose, as long as the rescaled masses remain within a numerically suitable range for the GP framework. All rescaled masses in higher-rank analytical expressions can be straightforwardly converted to actual masses in units of GeV or MeV using the most suitable relation from Table LABEL:tab:long6 or LABEL:tab:long7, followed by appropriate substitutions. For example, one can use $\delta=\exp(m^{\prime}_{\mu}/m^{\prime}_{\rho})$ from Table LABEL:tab:long7, where $m^{\prime}_{\mu}$ and $m^{\prime}_{\rho}$ denote the muon and $\rho(770)$ meson masses in GeV or MeV. From this expression, the value of $m^{\prime}_{\rho}$ can be determined and subsequently applied across all dimensionless expressions derived in this study. Determining the optimal analytic expression for the $\rho(770)$ mass, however, lies beyond the scope of this paper. This task should be pursued in future analyses of the derived analytic snippets, with the goal of identifying relationships among fundamental constants that enable the construction of physics models with a minimal set of parameters.
4 Conclusion
@@
In this paper, we derived analytic expressions - referred to as “snippets” - connecting the fundamental constants of the SM using an approach based on GP. We presented results up to analytic rank 10, limiting the solutions to a relative precision of at most 1%. The results with a relative difference greater than 1%, corresponding to the $u$- and $d$-quark masses, the CKM mixing angles, and the CP-violating phase, are not included in this article. Analytic expressions with ranks above 10, without any restriction on predicted precision $\Delta D$, and without rescaling by the $\rho(770)$ meson mass, are available [12] from the authors.
@@
A few known (though likely coincidental) relationships were successfully recovered. Surprisingly, the obtained GP results reveal a significant number of simple analytic patterns. Some of them can pass dimensional analysis. However, no evident patterns emerged from the relationships between the constants that could indicate an underlying simple mathematical structure with a limited number of free parameters. Although we expect that most of the expressions, undoubtedly, arise by pure chance, it is possible that some may hint at an underlying structure or deeper theory linking these fundamental constants - should such a theory exist. The coincidental nature of some (or all) of the snippets will be revealed over time as more precise data on the SM parameters become available.
@@
We believe that the compiled lookup library of analytic snippets will be useful for model builders. We also hope that the derived expressions will serve as a valuable resource for future AI-driven investigations, providing building blocks for uncovering a potential underlying law that connects all parameters of the SM through a small set of fundamental constants.

### v3

@@ 6 $\exp(x)$ and $\log(x)$ ($\log$ base of $e$ of $x$), $\cos(x)$, $\sin(x)$, $\tan(x)$, $\arctan(x)$
3.5 Validation of the GP Algorithm
As a sanity check, we have validated the genetic programming approach using known relationships. Although these relationships are not discussed in the context of the SM or any theory, they were established in the past, and have been deemed well suited for our purpose of testing the algorithm.
The first relationship is [1.3022455]:
@@ 3.5 Validation of the GP Algorithm
The second benchmark test uses the relation between the $Z$ boson mass ($m_{Z}$), the top-quark mass ($m_{t}$) and the Higgs boson mass ($m_{H}$) [Torrente-Lujan:2015jea]:
@@
Our program detected Equations (3.1) and (3.2) expressions using special test runs. The obtained simplicity ranks were 30 and 20, respectively. The difference $m_{H}-\sqrt{m_{Z}\>m_{t}}$ = 0.310-0.310 is greater than the absolute uncertainty allowed on the Higgs mass. When GP finds this equation, it is not considered a solution.
4 Results of the GP Algorithm
In this section, we discuss the output of the GP algorithm. In total, more than 1 million relations were generated up to simplicity rank 50 for each input, both with and without physical units. All of these relations are available at [GitRepo].
4.1 Output for Inputs with Physical Units
First, GP was used with the inputs from Table 1. Since the input includes a mix of variables with and without physical units, the vast majority of the resulting relations lack physical meaning. The dataset containing the GP outputs, along with markers indicating whether each expression satisfies dimensional analysis, is available in [GitRepo]. These markers were added by an external post-processing program that is not part of the GP framework. By applying this dimensional-analysis algorithm and visually inspecting expressions with a low analytic rank, we find that approximately 12% of the generated relations are dimensionally consistent.
Several of the simplest expressions that pass dimensional analysis will be discussed later.
4.2 Output for Dimensionless Inputs
We also generated GP snippets using Table 1, after replacing all masses with the dimensionless variables listed in Table 2. All expressions in the resulting GP output are dimensionally consistent, offering significant statistics and flexibility for future analyses without requiring additional dimensional analysis. Therefore, in this section, we provide several tables with such analytic snippets that demonstrate the GP output.
@@ 4.2 Output for Dimensionless Inputs
As an example, the expressions with the lowest analytic ranks, 6 and 7, are shown in Tables LABEL:tab:long6 and LABEL:tab:long7, respectively. The regularities appear rather trivial, and can be checked by hand. Table LABEL:tab:long10 of the Appendix Appendix shows the results with rank 10. Note that the analytic expressions with ranks 8 and 9 do not exist because of the default construction of the ranking system. We intentionally retained repeated snippets involving the same variables, as they may help model builders identify interesting patterns.
The tables also show the relative difference for each analytic snippet. It is defined as $\Delta D=100\%\times|y-y_{t}|/y_{t}$ (expressed as a percentage), where $y_{t}$ is the target value and $y$ is the predicted value. All relations have $\Delta D<\varepsilon^{rel}$ on the target constants according to the definition of GP. The distribution of the $\Delta D$ values will be discussed later. Only formulas with $\Delta D$ smaller than 1% are listed to reduce the sizes of the tables. This requirement truncates the presentation for some targets, which usually correspond to $u,d$-quark masses, the CKM mixing angles and CP-violating phase, since they have relative uncertainties $\varepsilon^{rel}$ above 1%; see Table 2. Other target constants should not be affected by the 1% requirement, since their $\Delta D$ values are smaller than $\varepsilon^{rel}$ listed in Table 1.
The results shown in Tables LABEL:tab:long6 and LABEL:tab:long10 are dominated by the expressions for $\delta$, which has the largest absolute and relative uncertainties ($\varepsilon$ and $\varepsilon^{rel}$) among all the input variables. Among all the ranks obtained, the lepton masses had the fewest solutions due to their relatively small mass uncertainty. All relationships from the symbolic regression have been preserved, even if they may appear highly unlikely. We remind that this choice reflects our intent to remain as theory-agnostic as possible, in case these analytic snippets are combined and simplified, and later used to inform or construct a model that connects all SM constants.
@@
We should recall that the introduction of the non-fundamental particle, the $\rho(770)$ meson, into our study was made to provide dimensionless inputs for GP, which are essential for the computational efficiency. The mass of this particle is an auxiliary parameter. It defines the limits of numerical precision for the GP and, consequently, the applicability of any potential theory connecting the SM constants, should such a theory exist. In principle, any particle can be used for this purpose, as long as the rescaled masses remain within a numerically suitable range for the GP framework. The analytic relations using the $\phi(1020)$-meson mass, $1019.461\pm 0.016$ MeV [ParticleDataGroup:2024cfk], are available from [GitRepo].
As a measure of how many expressions can be used for further study, Figure 1 shows the distribution of the values of $\Delta D$ (as defined above) for the GP outputs for the SM constants without scaling (filled histogram) and with the scaling by the $\rho$ and $\phi$ masses. Only expressions with an analytic rank of less than 40 are shown. A cutoff of 40 was chosen to avoid overly complex expressions.
As one can see from Figure 1, the number of GP expressions with scaled SM constants is significantly larger than the number of relations without mass scaling. We also recall that only 12% of the expressions without mass scaling pass dimensional analysis. A smaller number of GP expressions for the $\phi$-mass normalization, compared to the $\rho$-mass normalization, is due to the fact that the $\phi$-meson mass has a smaller relative experimental uncertainty than the $\rho$-meson mass.
Figure 1: The distribution of $\Delta D=100\%\times|y-y_{t}|/y_{t}$ (see the text) for the expressions without mass scaling (the filled histogram), and after scaling by the $\rho$ and $\phi$ meson masses. Only expressions with analytic rank smaller than 40 are shown.
Note that all rescaled masses in higher-rank analytical expressions can be straightforwardly converted to actual masses in units of GeV or MeV using the most suitable relation from Table LABEL:tab:long6 or Table LABEL:tab:long7, followed by appropriate substitutions. For example, one can use $\delta=\exp(m^{\prime}_{\mu}/m^{\prime}_{\rho})$ from Table LABEL:tab:long7, where $m^{\prime}_{\mu}$ and $m^{\prime}_{\rho}$ denote the muon and $\rho(770)$ meson masses in MeV. From this expression, the value of $m^{\prime}_{\rho}$ can be determined and subsequently applied across all dimensionless expressions derived in this study. This means that the dimensionless input allows for the generation of high-ranking, dimensionally consistent analytical expressions with relatively low CPU usage during GP production. Determining the optimal analytic expression for the $\rho(770)$ mass from the GP outputs, however, lies beyond the scope of this paper. This task should be pursued by model builders in future analyses of the derived analytic snippets, with the goal of identifying relationships among fundamental constants that enable the construction of physics models with a minimal set of parameters.
4.3 Discussion of the Results
In this section, we discuss the outputs generated by the GP algorithm that are common to both dimensional and dimensionless inputs. Outputs that satisfy dimensional analysis are more easily obtained when using dimensionless inputs, achieved by rescaling the masses. We recall that rescaling reduces the disparity in numerical values, particularly when considering lepton masses. The discussion below serves as an example of the step (2) described in Section 2. A complete set of requirements for this stage of the analysis should be addressed in a dedicated physics paper, which is beyond the scope of this work.
@@ 4.3 Discussion of the Results
A preliminary visual inspection of the lowest‑rank expressions reveals relationships that appear to be both remarkably simple and precise, warranting further investigation. For example, Table LABEL:tab:long10 has an expression $m_{t}=m_{W}+m_{W}\>\delta$ connecting the top-quark mass $m_{t}$ with the $W$ mass and $\delta$. Tables with analytic ranks above 10 are available from [GitRepo]. The rank 15 snippets listed in [GitRepo] contain another relation $m_{H}=m_{t}-(m_{b}+m_{t}/4)$ connecting the Higgs mass $m_{H}$ with $m_{t}$ and the $b$-quark mass. Another relation of the same rank is $m_{u}=2m_{e}/\sqrt{\theta_{12}}$, where $m_{u}$ is the mass of the $u$-quark, $m_{e}$ is the mass of the electron and $\theta_{12}$ is the CKM 12 mixing angle. The found expression $m_{H}=5\>m_{W}-3\>(m_{t}-m_{W})$ connects the Higgs mass with the top-quark mass and the $W$ mass. All such expressions are valid for the original masses with units, without scaling by the $\rho(770)$ mass. The precision of these analytic relations significantly exceeds that of Equation (3.2).
@@
We should mention that many relations involving the Higgs mass are valid only for the dimensionless constants listed in Table 2. For example, $m_{H}=m_{t}-6\>\sqrt{m_{W}}$ connects $m_{H}$, $m_{W}$ and $m_{t}$. It does not pass the dimensional analysis when using the original masses, i.e., without rescaling the masses by the $\rho(770)$ meson mass. However, the mass of this meson can easily be substituted using Table LABEL:tab:long6 or Table LABEL:tab:long7.
@@
To identify meaningful relationships from the GP output, it is necessary to apply constraints that reduce the number of relations arising from mere numerical coincidences. For the SM, such a constraint could be defined as follows: One of the known properties of the SM is that the masses of fermions, as well as the $m_{Z}$ and $m_{W}$ boson masses, tend toward zero when the Higgs boson mass is set to zero. This behavior arises from their dependence on the Higgs vacuum expectation value through the Yukawa couplings, which are free parameters in the theory [PhysRevLett.19.1264]. We identified several simplest relations connecting the quark and $Z/W$ masses with the Higgs mass, which have the above-mentioned property, and also pass the dimensional analysis. Here are some examples with small analytic ranks:
@@
These relations hold for the masses presented in Table 1 and for the rescaled masses in Table 2. Analytical expressions for lepton masses as linear functions of the Higgs mass were not found. Nevertheless, these examples illustrate that the reduction of the free parameters of the SM is possible, since 8 masses are expressed via the Higgs mass and 4 measured constants ($\alpha^{-1}$, $\delta$, $\theta_{12}$ and $\theta_{13}$). The origin of these expressions may be purely coincidental, and can only be understood within the framework of a specific model, which is beyond the scope of this paper.
@@
The obtained analytic expressions can be used to develop theories by model builders, or serve as inputs for various artificial intelligence techniques to discover hidden analytic structures. For example, the analytic snippets can be simplified or combined, variables can be substituted, and dimensional analysis can be applied after converting the masses back to their original physical units.
5 Next Steps
The goal of this paper is to explore the first step in the physics program: identifying the analytic structures among SM parameters. This objective is outlined in Section 2. Two approaches are discussed in this paper: (1) Using the original SM parameters with mixed units, followed by the application of post-processing algorithms to verify the dimensional correctness of the obtained expressions; (2) Generating dimensionally consistent expressions after normalizing by an auxiliary mass. The first approach produces smaller numerical noise (after dimensional analysis), while the second approach results in a significant number of dimensionally correct expressions, offering greater flexibility for model builders. However, in the latter case, the challenge shifts from the GP itself to the analytic substitutions necessary to remove the dependence on the auxiliary mass. It should be noted that there is significant overlap between these two approaches, since dimensionally correct expressions from method (1) can also be found among the expressions of method (2). We expect that the numerical noise is higher in the second approach.
On the technical side, several improvements can be made for GP, such as incorporating variations of arguments in the discovered expressions and performing dimensional checks inside the GP algorithm. The latter may improve the efficiency of GP, since external algorithms for checking dimensional correctness would no longer be required.
The inclusion of cosmological constants is another possible extension, which can be implemented in a straightforward manner, since the analysis code is publicly available. However, one should keep in mind the limitation of the Picat program: it cannot handle values smaller than $10^{-17}$. For this reason, this paper did not use the Planck mass, which is more natural in particle physics for removing mass units.
Even when dimensionally consistent expressions are identified, this does not completely eliminate the numerical noise arising from coincidental relations. The second step, as discussed in Section 2, must therefore involve filtering out coincidental expressions with the goal of identifying relations that can be given a physical interpretation within some unification model. Although this step is beyond the scope of the present paper, here, we should mention one statistical method that can be used to address this problem. Recently, an approach called the “random sampling test” was introduced [Chekanov:2025pmb], in which random smearing is applied to the GP inputs followed by re-evaluation of the simplicity ranks. If a consistent system of the relations exhibits the lowest simplicity rank among all random variations, it may suggest a genuine interdependence among the fundamental constants within a high-dimensional functional space. This method was found to be CPU-intensive, since every variation of the GP inputs requires the creation of new analytic snippets, followed by dimensional analysis, and re-calculation of simplicity ranks. To address high CPU usage in the future, large-scale high-performance computing resources may be required.
It is important to emphasize that any meaningful physical model or theory must be falsifiable. This means that as measurements of SM constants become more precise, the GP relations used by such a model or theory can be retested; if they fail to hold, that mathematical framework is not viable. In this sense, the numerical noise problem is itself falsifiable.
6 Conclusions
This paper discusses a promising method based on genetic programming for uncovering underlying mathematical relationships directly from data. The first step follows a symbolic regression approach: deriving analytical expressions from the data and generating a large set of candidate relations. The second step involves identifying simplest connecting patterns among these relations, dominated by numerical coincidences. In this paper, this step leverages dimensional analysis and general SM expectations to reveal possible analytic structures within a high-dimensional functional space. To generalize the use of this method, the SM constraints can be substituted with any known conditions pertinent to the expected outcome, allowing for the filtering of expressions that are likely artifacts of numerical noise.
@@ 6 Conclusions
The derived analytic expressions—referred to as “snippets”—connect the fundamental constants of the SM using an approach based on GP. We presented results up to analytic rank 10, limiting the solutions to a relative precision of at most 1%. The results with a relative difference greater than 1%, corresponding to the $u$- and $d$-quark masses, the CKM mixing angles, and the CP-violating phase, are not included in this article. Analytic expressions with ranks above 10, without any restriction on predicted precision $\Delta D$, and with and without rescaling by the $\rho(770)$ (or $\phi(1020)$) meson mass, are available [GitRepo] from the authors. The same repository includes a basic example with the analysis code and the configuration files for the Picat program [PicatWebsite].
@@
A few known (though likely coincidental) relationships were successfully recovered. Surprisingly, the obtained GP results reveal a significant number of simple analytic patterns. Some of them can pass dimensional analysis. However, no evident patterns emerged from the relationships between the constants that could indicate an underlying simple mathematical structure with a limited number of free parameters. Although we expect that most of the expressions, undoubtedly, arise by pure chance, it is possible that some may hint at an underlying structure or deeper theory linking these fundamental constants—should such a theory exist. The coincidental nature of some (or all) of the snippets will be revealed over time as more precise data on the SM parameters become available.
@@
We believe that the compiled lookup library of analytic snippets will be useful for model builders. We also hope that the derived expressions will serve as a valuable resource for future AI-driven investigations, providing building blocks for uncovering a potential underlying law that connects all parameters of the SM through a small set of fundamental constants. Finally, the proposed method can be extended to other research fields characterized by numerous free parameters, such as the study of cosmological constants and various branches of applied physics.